OpenStudy (anonymous):

find the limit of 3x^2+6x-9/2x^2-2 as x approaches 1

OpenStudy (asnaseer):

You might find this site useful: http://archives.math.utk.edu/visual.calculus/1/limits.7/ It has good walk-through tutorials on this topic.

OpenStudy (anonymous):

you basically just need to brush up on factoring

OpenStudy (anonymous):

(a+b)^2=a^2+2ab+b^2, if you factor a 3 out of that numerator, you will have this situation

OpenStudy (anonymous):

yeah how do i factor 9 to add up to 6

OpenStudy (anonymous):

or do i not have to

OpenStudy (anonymous):

factor out a 3 first, it will make it much easier

OpenStudy (anonymous):

ive got (x+3)(x-3)

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

thats a difference of 2 squares

OpenStudy (anonymous):

so what do i do

OpenStudy (anonymous):

if you expanded that you'd get x^2-9

OpenStudy (anonymous):

this part 3x^2+6x-9

OpenStudy (anonymous):

factor out a 3

OpenStudy (anonymous):

3,6,9 all have a common factor of 3 here

OpenStudy (anonymous):

3(x^2 +6x-9)

OpenStudy (anonymous):

you have to factor the 3 out of the whole numerator so that when you multiply it you get the original expression back

OpenStudy (anonymous):

what you just did will yield 3x^2 + 18x -27

OpenStudy (anonymous):

3(x^2-2x-3)

OpenStudy (anonymous):

+2x

OpenStudy (anonymous):

good

OpenStudy (anonymous):

now factor the denominator

OpenStudy (anonymous):

very much the same thing, guess which number you're pulling out?

OpenStudy (anonymous):

2(x^2-1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now factor (x^2-1)

OpenStudy (anonymous):

(difference of 2 squares), i mentioned it a little bit ago

OpenStudy (anonymous):

hmm im not sure here

OpenStudy (anonymous):

hehe you just did it mistakenly earlier

OpenStudy (anonymous):

remember you typed (x-3)(x+3)

OpenStudy (anonymous):

oh that

OpenStudy (anonymous):

(x+1)(x-1)

OpenStudy (anonymous):

yes, now it's all factored. put it all together and cancel like terms

OpenStudy (anonymous):

ok but could i have plug in 1 and get an answer

OpenStudy (anonymous):

actually, you can further factor the numerator

OpenStudy (anonymous):

right now you have the numerator at (x^2+2x-3)

OpenStudy (anonymous):

you could not have p lugged 1 into the original equation before factoring it

OpenStudy (anonymous):

look at the original denominator

OpenStudy (anonymous):

2x^2-2

OpenStudy (anonymous):

if you plugged a 1 in here, you'd get a divide by 0 error

OpenStudy (anonymous):

i see you will just get 0/0

OpenStudy (anonymous):

so try factoring x^2+2x-3

OpenStudy (anonymous):

(x-1)(x+3)

OpenStudy (anonymous):

good

OpenStudy (anonymous):

now everything is factored

OpenStudy (anonymous):

so when i cancel i get x+3/x+1

OpenStudy (anonymous):

don't forget about the 3 and 2 you factored out

OpenStudy (anonymous):

3(x+3)/(2(x+1)

OpenStudy (anonymous):

at this point you can plug in 1 and you will not get a divide by 0 error

OpenStudy (anonymous):

because you cancelled out the (x-1)

OpenStudy (anonymous):

ok i get 12/4

OpenStudy (anonymous):

when i reduce i get 3/1

OpenStudy (anonymous):

yes. that's it

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

3

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i have another limit problem but its approaching infinity

OpenStudy (anonymous):

ok, let's have it

OpenStudy (anonymous):

x^2 + 2x-3/2x^2-2 as x approaches infinity, what is the limit

OpenStudy (anonymous):

now look at this numerator and denominator

OpenStudy (anonymous):

you already factored them in the last problem

OpenStudy (anonymous):

give me your factored equation for this problem

OpenStudy (anonymous):

(x-1)(x+3)/2(x^2-1)

OpenStudy (anonymous):

the denominator further factors out to (x-1)(x+1)

OpenStudy (anonymous):

yes, now cancel

OpenStudy (anonymous):

x+1/x+1

OpenStudy (anonymous):

or 2x+1

OpenStudy (anonymous):

what happened to x+3?

OpenStudy (anonymous):

my final looks like this (x+3)/(2x+1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x+2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2(x+1) then you foil

OpenStudy (anonymous):

that's as far as you can factor it: (x+3)/[2(x+1)]

OpenStudy (anonymous):

you cant do 2x+2?

OpenStudy (anonymous):

what's your thinking on that? just trying to see where you're going wrong

OpenStudy (anonymous):

oh my mistake, i see what you were doing

OpenStudy (anonymous):

thats not a big deal to expand the denominator to 2x+2

OpenStudy (anonymous):

so youd have (x+3)/(2x+2)

OpenStudy (anonymous):

i thought you meant the whole equation factored out to that

OpenStudy (anonymous):

if you have 2(x+1)...why not just foil (2)(x)=2x...2(1)=2...2x+2...

OpenStudy (anonymous):

you can, but it's generally not necessary after you've already factored

OpenStudy (anonymous):

plus if you do it too early you'll have to factor it again later

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so since i can't plug in infinity...what should i do

OpenStudy (anonymous):

now you have to think what would happen to f(x) is you kept plugging in very large numbers

OpenStudy (anonymous):

it would remain large

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

try plugging in 100 just to test it

OpenStudy (anonymous):

well it might cut down...but at the end it would still be large

OpenStudy (anonymous):

100+3/2(100+1)=103/202= .50

OpenStudy (anonymous):

ok now 1000

OpenStudy (anonymous):

the same .50

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 1/2?

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (anonymous):

sure